# Sweep Angle Effects of Flow Over an Undulated Cylinder

### v1

@@ Abstract
Flow over a seal whisker-inspired undulated cylinder at swept back angles is computationally investigated, comparing the vortex shedding, forces, and wake characteristics to those of an equivalent smooth geometry. Numerous prior studies have demonstrated that undulated cylinders can reduce mean drag and unsteady lift oscillations; however, none have isolated the effects of sweep resulting from whisker positioning in flow. Relevant not only to understanding seal whiskers, such findings are also useful in engineering structural applications to reduce drag and vortex-induced vibration. This investigation performs direct numerical simulation of flow across a rigid, infinite-span, undulated cylinder at sweep angles from 0 to 60 and at Reynolds numbers of 250 and 500. At zero sweep, the undulated cylinder breaks up coherent two-dimensional vortices, having the effect of reducing drag by 10.4% and root mean square lift by 91.2% compared to a smooth elliptical cylinder. With sweep added, the prominence of spanwise vortex breakup and force suppression is reduced, approximating flow over smooth ellipse geometry as sweep increases. At low sweep angles of 15 and 30 degrees, lift is still suppressed by 75.5% and 50% while drag results in a smaller difference of 7.4 and 1.9% reduction from a smooth ellipse.
@@ 1 Introduction
The flow physics and force reduction properties of swept flow over a seal whisker-inspired undulated cylinder geometry are explored. Distinct from the smooth whiskers of other mammals, most species of seal have undulated whiskers that consist of two sets of opposing undulations pictured in Figure 1. The undulated whisker shape enhances a seals’ excellent navigation and prey tracking abilities enabled through hydrodynamic sensing [1, 2, 3]. This capability comes from the observation that seal whiskers significantly reduce drag and oscillating lift forces compared to smooth whiskers, improving their effective signal to noise ratio when sensing their environment [4, 5]. Despite a rich background in research [6, 7] that includes whisker anatomy and sensing abilities [1, 2, 3], as well as investigations into how whisker shape produces unique flow effects [8, 9, 10, 11, 5, 12, 13, 14, 15, 16, 17, 4, 18, 19, 20, 21, 22, 3, 23, 24], the influence of sweep angle, defined as the angle with which whiskers are swept back relative to the freestream flow relative to their perpendicular orientation before sweep, is yet to be investigated. An important overlapping field of study is the flow at swept (or yaw) angles over bluff-bodies, which is of interest in multiple disciplines from aerodynamics to civil engineering [25, 26, 27, 28, 29, 30, 31, 32, 33].
A common feature in research of seal whisker-inspired geometry is the variation in—and impact of—geometric parameters of the surface. A description commonly used in fluid dynamics investigations is the whisker model introduced by Hanke et al. [4] which was based on measurements of excised harbor seal whiskers. Defined by two angled ellipse cross sections, a seven variable parametric description of the three-dimensional surface was created. From this, Lyons et al. [5] redefined the model into independent variables and established a variable ranking based on hydrodynamic relevance [5]. The geometric variables that produced the most significant impact on forces, shown in Figure 1, were wavelength, aspect ratio, and undulation amplitudes. Other studies performed analysis focusing on specific geometric parameters such as undulation wavelength to probe its impact on vortex shedding, forces, and shedding frequency [13, 24, 34, 35]. These studies established the capabilities of the whisker-inspired geometry which may be extended to applications beyond those found in the original biological system.
Whiskers in the field experience a wide range of flow directions, speeds, and non-uniformity in their tumultuous flow environment. Angle of attack effects (rotation about the $\mathit{z}$-axis) have been investigated previously for span-perpendicular flow, finding that reduced forces are most prevalent at zero angle of attack while the flow response increasingly approximates flow over a bluff body as the angle of attack reaches 90 degrees [20, 36, 10]. Whiskers also encounter variation in incoming flow angle due in equal parts to their natural curvature, bending in response to incoming flow, and due to a seals’ ability to influence their orientation via their facial musculature. Thus, this research explores the hydrodynamics properties of flow at oblique angles to the whisker span, defined commonly as swept angle in aerodynamics or yaw angle in bluff bodies.
Sweep angle, defined by $\mathit{\Lambda}$ in Figure 1, where $\mathit{\Lambda}=0$ corresponds to the perpendicular flow investigated in prior research, is comparable to research of bluff body flow behavior with a spanwise velocity component. Previous exploration of swept flow drew comparisons to the ubiquitous problem of crossflow over a circular cylinder and include experiments [25, 26, 27, 28, 29], often dependent on finite aspect ratio and end effects, and simulations [30, 31, 32, 29, 33] over a range of Reynolds number regimes. These studies often test the "crossflow principle" or "independence principle" [37] which estimates the effects of sweep on fluid forces and shedding characteristics by scaling perpendicular flow values proportionally to the normal component of flow over a swept body. This idea does not take into account three-dimensional flow effects, but can still give reasonable estimates of forces for rigid cylinders with minimal end condition effects, and for a range of sweep angles generally limited to less than or equal to 45 degrees [30, 26, 25, 32, 28, 31].
Through simulation of flow over an infinite-span whisker model within a range of sweep angles from 0 to 60 degrees, this paper investigates how sweep impacts the force-reduction properties of an undulated cylinder. The following sections are organized into Section 2 where the whisker geometry, numerical model, and flow parameters are introduced, Section 3 in which results are summarized, and Section 4 which contains concluding remarks.
@@
Figure 1: Geometric variables comprising the undulated cylinder model as defined by Lyons et al. [5].
@@ 2 Methods
${\partial t}+\textbf{u}\cdot\nabla\textbf{u}=-\nabla p+\frac{1}{Re}\nabla^{2}\textbf{u}$ u (1)
@@
The numerical solution is discretized using OpenFOAM [38] second-order finite volume libraries, and solved with a pressure-implicit split-operator (PISO) algorithm. The numerical method follows that of prior work [34, 13, 35].
@@
The undulated geometry is comprised of two sets of undulations, one in the chord ($x$-direction), and another in the thickness ($y$-direction), as seen in Figure 1. The surface undulations are prescribed from the hydrodynamically relevant non-dimensional parameters developed by Lyons et al. [5] with mean thickness $T=1$, aspect ratio $\gamma=C/T=1.919$, wavelength $\lambda/T=3.434$, transverse amplitude $A_{T}/T=0.09$, streamwise amplitude $A_{C}/T=0.23$, undulation offset $\epsilon=0.34$, and undulation symmetry $\phi=0.015$. These values yield a surface with the same measurements as that originally defined by Hanke et al. [4], and the conversion between the two geometric definitions is provided in Lyons et al.
@@
The analytical surface of the whisker is generated and defined using the mesh morphing algorithm described by Yuasa et al. [13]. Starting from a circular cylinder computational domain with a radially structured mesh, the algorithm radially translates the surface mesh nodes to morph the cylinder into a whisker geometry with the specified geometric features defined above. The result is a body-fitted, structured mesh of an undulated cylinder that is repeatable and independent of changes to the starting cylinder mesh resolution.
@@
The domain is held constant through a range of swept angle flows which are created by an angled inlet velocity at angle $\mathit{\Lambda}$ with $\mathit{\Lambda}=0$ corresponding to flow normal to the cylinder span. Due to the phase difference in leading and trailing edge undulations, the geometry possesses a directional bias relative to the sign of sweep angle. The positive $\mathit{\Lambda}$ direction is chosen such that angled flow is in line with the angled ellipses that define the whisker surface as visible in Figure 1. In the current investigation, only the positive (aligned with the undulation offset $\epsilon$) range of sweep angle $\mathit{\Lambda}$ is explored.
The cylindrical structured mesh shown in Figure 2(a) consists of 170 cells in the spanwise direction resulting in a spanwise resolution of $\Delta z/T=0.041$, 160 about the azimuthal direction, and 154 in the radial direction which are initially clustered to produce a minimum cell thickness of $\Delta r/T=0.003$ and radial resolution of $\Delta\theta=4.01^{\circ}$. While the mesh and computational domain have been previously validated at zero sweep angle [13, 34], a mesh independence study at sweep angle $\mathit{\Lambda}=60$ is documented in Section A.
Simulations are performed at Reynolds numbers of 250 and 500, which are biologically relevant based on seal swimming speeds and average whisker diameter [39]. Simulation time is specified in non-dimensional convective time units (CTU), or the time it takes for the flow to move one mean thickness at freestream velocity. All simulations begin with an initial transient time of 100 CTU to allow ample time for the flow to fully develop. After this point, a subsequent 400 CTU are simulated and utilized for all averaged quantities.
Forces due to pressure and skin friction are computed along the whisker surface as a function of time, and are used to compute the lift and drag coefficients. As sweep angle rotation occurs in the $xz$ plane, the magnitude of lift forces oscillating parallel to the $y$-axis require no rotation, whereas drag force will shift directions with changes in $\mathit{\Lambda}$ and is thus defined parallel to the freestream flow. The drag and lift coefficients are given by
$C_{D}=\frac{2F_{D}}{\rho U_{\infty}^{2}TL_{z}}\quad\quad\quad C_{L}=\frac{2F_{L}}{\rho U_{\infty}^{2}CL_{z}}$ (4)
where drag force is normalized by the whisker mean thickness $T$ (frontal area in the $yz$ plane), and lift is normalized by the mean chord length $C$ (an average area in $xz$). As these forces constantly oscillate in response to vortex shedding, force comparison between orientations is facilitated by mean values of drag. As the whisker geometry lacks camber and is symmetric about the $xz$ plane, lift forces oscillate about zero. The magnitude of oscillations is captured by the root mean square (RMS) of the coefficient over $n$ samples in time with a sampling rate of 50 samples per convective time unit,
$C_{L,RMS}=\sqrt{\frac{1}{n}\sum_{i=1}C_{L_{i}}^{2}}$ (5)
To provide a baseline comparison for forces measured from whisker-inspired undulated cylinder geometry, simulations of flow over a circular cylinder and elliptical cylinder are also performed. The circular cylinder maintains the same mean thickness (diameter) as the undulated geometry while the ellipse has the same mean aspect ratio. The elliptical cylinder captures the effect of a streamlined cross-section and serves as a useful intermediate comparison between a circular cylinder and the undulated cylinder.
For flow visualization, $Q$-criterion is calculated to identify vortices in the wake by relating the magnitude of local strain rate $\textbf{{S}}=\frac{1}{2}(\nabla\textbf{{u}}+\nabla\textbf{{u}}^{T})$ and rotation rate $\bm{\varOmega}=\frac{1}{2}(\nabla\textbf{{u}}-\nabla\textbf{{u}}^{T})$ tensors in the equation,
$Q=\frac{1}{2}(||\bm{\varOmega}||^{2}-||\textbf{{S}}||^{2}).$ (6)
Lastly, the distribution of energy in the wake is compared via contours of turbulent kinetic energy (TKE), defined by the half-sum of variances $\overline{(\textbf{{u}}^{\prime})^{2}}$ in fluctuating velocity $\textbf{{u}}^{\prime}$ obtained by subtracting time averaged velocity $\overline{\textbf{{u}}}$ from instantaneous velocity $\textbf{{u}}(t)$.
$k=\frac{1}{2}\left[\overline{(u^{\prime})^{2}}+\overline{(v^{\prime})^{2}}+\overline{(w^{\prime})^{2}}\right]$ (7)

### v2

# Sweep Angle Effects of Flow Over a Seal Whisker-Inspired Undulated Cylinder

@@ Abstract
Flow over a seal whisker-inspired undulated cylinder at swept back angles is computationally investigated, comparing the vortex shedding, forces, and wake characteristics to those of an equivalent smooth geometry. Numerous prior studies have demonstrated that undulated cylinders can reduce mean drag and unsteady lift oscillations; however, none have isolated the effects of the sweep angle resulting from whisker positioning in flow. Inspired by the active control seals exert over their whiskers while navigating and sensing in unsteady aquatic environments, this study investigates how such orientation influences the hydrodynamic performance of the geometry.
Simulations are performed of flow across a rigid, infinite-span, undulated cylinder at sweep angles from 0 to 60 and at Reynolds numbers of 250 and 500. At zero sweep, the undulated cylinder breaks up coherent two-dimensional vortices, having the effect of reducing drag by 11.4% and root mean square lift by 90.8% compared to a smooth elliptical cylinder. With sweep added, the prominence of spanwise vortex breakup and force suppression is reduced, approximating flow over smooth ellipse geometry as sweep increases. At low sweep angles of 15 and 30 degrees, lift is still suppressed by 72.4% and 47.6% while drag results in a smaller difference of 5.7 and 1.6% reduction from a smooth ellipse. These results reinforce that sweep angle is a significant parameter both mechanically and biologically in the flow physics of whisker-inspired undulated geometries.
@@ 1 Introduction
Distinct from the smooth whiskers of other mammals, most species of true seal (such as the harbor seal pictured in Figure 1) have undulated whiskers [morgenthal_characterization_2025]. The whisker shape, comprised of two opposing sets of undulations as depicted in Figure 2, enhances seals’ excellent navigation and prey tracking abilities enabled through hydrodynamic sensing [dehnhardt_seal_1998, dehnhardt2001, Murphy2017]. The undulated geometry reduces drag and oscillating lift forces compared to smooth whiskers, improving the effective signal to noise ratio when sensing their environment [hanke_harbor_2010, lyons_flow_2020].
Prior research on seal whiskers began primarily with biological and behavioral studies which established that undulated whiskers play a central role in hydrodynamic sensing, enabling seals to detect water motions and track wakes generated by prey [dehnhardt_seal_1998, dehnhardt2001, Murphy2017]. These findings motivated extensive investigation into the physical mechanisms underlying this geometry’s modification of flow compared to canonical bluff bodies [zheng_review_2021]. To address these questions, experimental and numerical studies demonstrated that the undulated whisker geometry operates by suppressing vortex-induced vibrations and reduces mean drag and unsteady lift force compared to smooth cylinders by disrupting spanwise vortex coherence [hanke_harbor_2010, witte_wake_2012, wang_wake_2016, liu_phase-difference_2019]. Parametric modeling studies were then undertaken to isolate the geometric features most responsible for these effects, identifying wavelength, aspect ratio, curvature, and undulation amplitude as dominant contributors to force reduction and wake modification [lyons_flow_2020, yuasa_simulations_2021, lyons_dye_2021, kamat_wavelength_diameter_2024, dunt_frequency_2024, yuanji_curvature_2025].
In parallel with these hydrodynamic investigations, a growing body of work has examined how altered wake structure and force response contribute to the whisker’s sensing function as a biological structure or man-made bioinspired sensor. Experimental and computational studies have demonstrated the ability of whisker-inspired structures to detect and track upstream wakes, often incorporating fluid–structure interaction or imposed upstream disturbances to model realistic sensing scenarios, where whisker motion or environmental disturbances are central to the problem formulation [beem_calibration_2013, beem_wake-induced_2015, kamat_wavelength_diameter_2024, zheng_printed_whisker_sensor_2022, eberhardt_whisker_sensor_2016, geng_whisker_sensor_2025, liu_phase-difference_2019, zhu_FIV_tandem_whiskers_2026, liu_flow-signal-correlation_2022]. Beyond biological analogs, related flow phenomena have also been observed for non-biological wavy or sinusoidal cylinders, where spanwise geometric modulation disrupts vortex shedding and alters force response [ahmed_transverse_1992, lam_effects_2009, lam_three-dimensional_2004, zhang_piv_2005]. However, such idealized geometries differ substantially from seal whiskers in cross-sectional shape, undulation pattern, and biological scaling.
Despite this extensive background, seal whisker research has almost exclusively considered span-perpendicular inflow. While angle of attack effects under perpendicular flow have been examined [murphy_effect_2013, kim_effect_2017], the influence of sweep angle, where the flow is no longer perpendicular to the whisker axis, has not been systematically investigated. This gap is particularly relevant given the active and variable orientation of whiskers during swimming and foraging.
As shown by Figure 1, seals exert active musculature control to protract their whiskers [milne_orient_control_2020, adachi_prey_sensors_2022]. In particular, during active foraging, elephant seals have been recorded alternating between protracting their whiskers outward into flow, particularly when prey is present, and sweeping them back against their face [adachi_prey_sensors_2022]. Additionally, individual whiskers exhibit intrinsic curvature [luo_intrinsic_curvature_2023, rinehart_characterization_2017] and are distributed across the face with varying spatial orientation [dougill_whisker_morphology_2023, graff_spatial_arrangement_2024]. Combined with head and neck motion during swimming, these factors introduce natural variation in flow direction relative to each whisker, making sweep angle a relevant variable in this system. Despite this biological relevance, the effects of sweep on seal whiskers have not been previously studied in detail, as it is difficult to capture nuanced whisker data from live seals or excised whiskers [Murphy2017]. Computational simulations, however, provide a valuable tool for isolating specific geometric and orientation variables in seal whisker hydrodynamics that are difficult to capture in experimental measurements or biological samples. In addition, computational studies of whisker-inspired geometries can inform engineering applications, such as flow sensing, drag reduction, and wake control, where analogous three-dimensional undulations or orientation effects may be exploited.
(a) Whiskers in swept-back, relaxed position
(b) Whiskers protracted away from face
Figure 1: Harbor seal Sprouts, UCSC Pinniped Lab demonstrating a range of whisker movement. (National Marine Fisheries Service (NMFS) permit 14535) Panel (b) adapted from murphy_navy_seals_2021.
A three-dimensional surface model of a seal whisker was introduced by hanke_harbor_2010, and defined by two angled ellipse cross sections and a distance between them. lyons_flow_2020 redefined the model into independent and hydrodynamically relevant variables and established the sensitivity of each parameter, finding wavelength, aspect ratio, and undulation amplitudes to be most impactful on forces. Follow-up studies performed analysis focusing on specific geometric parameters[yuasa_simulations_2021, lyons_dye_2021, lyons_wavelength_2022, dunt_frequency_2024, kamat_wavelength_diameter_2024, yuanji_curvature_2025], documenting the capabilities of the whisker-inspired geometry such that they may be extended to bio-inspired engineering applications. Angle of attack effects (rotation about the $\mathit{z}$-axis) have been investigated previously for span-perpendicular flow, finding that reduced forces are most prevalent at zero angle of attack while the flow response increasingly approximates flow over a bluff body as the angle of attack reaches 90 degrees [murphy_effect_2013, wang_flow_2017, kim_effect_2017, geng_angle-of-attack_VIV_2024].
Here, sweep angle is defined by $\mathit{\Lambda}$ in Figure 2, where $\mathit{\Lambda}=0$ corresponds to pure perpendicular flow. Commonly called sweep angle in aerodynamics, alternate terminology such as yaw angle or crossflow are often utilized in bluff body hydrodynamics. Research of crossflow over a smooth circular cylinder include experiments [king_vortex_1977, ramberg_effects_1983, williamson_oblique_1989, snarski_flow_2004, thakur_wake_2004] and simulations [wang_effect_2019, lucor_effects_2003, marshall_wake_2003, thakur_wake_2004, zhao_direct_2009] over a range of Reynolds number regimes. These studies serve as a reference for the expected behavior of bluff bodies at sweep angles, and represent a useful baseline for understanding the comparatively unique flow over the undulated cylinder. These works often invoke the classical independence principle (IP), which estimates forces on a swept body by scaling perpendicular flow values with the normal velocity component. While IP neglects three-dimensional flow effects, it can reasonably reproduce projected-area scaling and associated mean drag trends for some bluff-body geometries at moderate sweep angles. However, IP does not account for changes in wake instability, vortex shedding coherence, or spanwise phase variation, and therefore becomes less reliable for predicting forces by vortex shedding dynamics [ramberg_effects_1983].
This paper presents simulations of flow over an infinite-span model with sweep angles from 0 to 60 degrees, investigating how sweep impacts the force-reduction properties of an undulated cylinder. The sweep angles considered here are not intended to represent specific biological postures, but rather span a representative range of effective whisker orientations arising from curvature, placement, and motion. The intent is to isolate how sweep modifies force suppression and wake structure, rather than to prescribe a particular behavioral configuration. Section 2 introduces the geometry, the numerical model, and the flow parameters, whereas Section 3 presents results and discussion, and concluding remarks are summarized in Section 4.
@@
Figure 2: Geometric variables comprising the undulated cylinder model as defined by lyons_flow_2020.
@@ 2 Methods
$\frac{\partial\textbf{u}}{\partial t}+\textbf{u}\cdot\nabla\textbf{u}=-\nabla p+\frac{1}{Re}\nabla^{2}\textbf{u}$ (1)
@@
The numerical solution is discretized using OpenFOAM [openfoam] second-order finite volume libraries, and solved with a pressure-implicit split-operator (PISO) algorithm. The numerical method follows that of prior work [lyons_wavelength_2022, yuasa_simulations_2021, dunt_frequency_2024].
@@
The undulated geometry is comprised of two sets of undulations, one in the chord ($x$-direction), and another in the thickness ($y$-direction), as seen in Figure 2. The surface undulations are prescribed from the hydrodynamically relevant non-dimensional parameters developed by lyons_flow_2020 with mean thickness $T=1$, aspect ratio $\gamma=C/T=1.919$, wavelength $\lambda/T=3.434$, transverse amplitude $A_{T}/T=0.09$, streamwise amplitude $A_{C}/T=0.23$, undulation offset $\epsilon=0.34$, and undulation symmetry $\phi=0.015$. These values yield a surface with the same measurements as that originally defined by Hanke et al. which enables comparison to prior results of the same geometry at a similar Reynolds number. The conversion between the two geometric definitions is provided in lyons_flow_2020.
@@
The analytical surface of the whisker is generated and defined using the mesh morphing algorithm described by yuasa_simulations_2021. Starting from a circular cylinder computational domain with a radially structured mesh, the algorithm radially translates the surface mesh nodes to morph the cylinder into a whisker geometry with the specified geometric features defined above. The result is a body-fitted, structured mesh of an undulated cylinder that is repeatable and independent of changes to the starting cylinder mesh resolution.
The computational domain, shown in Figure 3(a), consists of an outer cylinder with a radius of 75TT, with a body-fixed computational coordinate system defined such that $x$ and $y$ are perpendicular and $x$ is aligned to the major (chord) axis, $y$ with the shorter thickness axis, and $z$ is in the spanwise direction. The outer boundary is divided into an upstream inlet with uniform velocity and zero-gradient pressure and a downstream, zero-gradient velocity and fixed pressure outlet. The model span consists of two wavelengths of the undulated cylinder identified in prior literature as sufficient to resolve the periodic flow structures created [kim_effect_2017], with periodic boundary conditions in the $z$-direction. Because the present spanwise domain length was selected based on resolution requirements for perpendicular inflow, and no span-length sensitivity study was conducted for strongly swept configurations, the possibility of longer-wavelength instabilities unique to high sweep cannot be fully excluded. However, the observed wake structures and force trends do not exhibit signatures of unresolved large-scale spanwise modes.
Sweep angles are varied by prescribing an angled inlet velocity relative to the body-fixed coordinate system. For the purpose of force decomposition, a rotated, freestream-aligned coordinate system ($x^{\prime}$, $y^{\prime}$, $z^{\prime}$) is introduced where $x^{\prime}$ is in the freestream direction, $^{\prime}=yy$, and $z^{\prime}$ is perpendicular to the ($x^{\prime}$, $y^{\prime}$) plane. Due to the phase difference in leading and trailing edge undulations, the geometry possesses a directional bias relative to the sign of sweep angle. The positive $\mathit{\Lambda}$ direction is chosen such that angled flow is in line with the angled ellipses that define the whisker surface as visible in Figure 2. In the current investigation, only the positive (aligned with the undulation offset $\epsilon$) range of sweep angle $\mathit{\Lambda}$ is explored.
The cylindrical structured mesh shown in Figure 3(a) consists of 170 cells in the spanwise direction resulting in a spanwise resolution of $\Delta z/T=0.041$, 160 in the azimuthal direction, and 154 in the radial direction which are initially clustered to produce a minimum cell thickness of $\Delta r/T=0.003$ and radial resolution of $\Delta\theta=4.01^{\circ}$. While the mesh and computational domain have been previously validated at zero sweep angle [yuasa_simulations_2021, lyons_wavelength_2022], a mesh independence study at sweep angle $\mathit{\Lambda}=60$ is documented in Section A.
Simulations are performed at Reynolds numbers of 250 and 500, which are biologically relevant based on seal swimming speeds and average whisker diameter [Lesage_classification_1999]. Both values lie within the same low-to-moderate Reynolds number regime commonly studied in whisker-inspired bluff-body flows and have been previously observed to demonstrate two distinct shedding modes from the same geometry [lyons_flow_2020, lyons_dye_2021, yoon_effect_2020, lam_effects_2009, hanke_harbor_2010, lyons_wavelength_2022, yuasa_simulations_2021, dunt_frequency_2024, kim_effect_2017, morrison_simulating_2016]. By examining two Reynolds numbers, the robustness of sweep-induced changes in force suppression and wake structures is explored over these two regimes. Prior simulations show that $Re=250$ produces clean and repeatable hairpin vortices, whereas at $Re=500$, these structures begin to break up and transition in the near wake [lyons_flow_2020, lyons_wavelength_2022, yuasa_simulations_2021, dunt_frequency_2024].Simulation time is specified in non-dimensional convective time units (CTU), or the time it takes for the flow to move one mean thickness at freestream velocity. All simulations begin with an initial transient time of 100 CTU to allow ample time for the flow to fully develop. After this point, a subsequent 400 CTU are simulated and utilized for all averaged quantities.
Forces due to pressure and skin friction are computed along the whisker surface as a function of time, and are used to compute the lift and drag coefficients. As sweep angle rotation occurs in the $xz$ plane, the magnitude of lift forces oscillating parallel to the $y$-axis require no rotation, whereas drag force will shift directions with changes in $\mathit{\Lambda}$ and is thus defined parallel to the freestream flow along $x^{\prime}$. The drag and lift coefficients are given by
$C_{D}=\frac{2F_{D}}{\rho U_{\infty}^{2}TL_{z}}\quad\text{and}\quad C_{L}=\frac{2F_{L}}{\rho U_{\infty}^{2}CL_{z}}$ (4)
where drag force is normalized by the whisker mean thickness $T$ (frontal area in the $yz$ plane), and lift is normalized by the mean chord length $C$ (an average area in $xz$). As freestream velocity magnitude is held constant, increasing sweep angle reduces the velocity component perpendicular to the span which is often responsible for driving dominant forces. As a result, a systematic reduction in force magnitudes occurs when normalizing coefficients with $U_{\infty}$. To aid in examination of force trends across sweep angle, an alternative normalization based on the perpendicular velocity component $U_{\perp}=U_{\infty}cos(\mathit{\Lambda})$ is
$C_{D\perp}=\frac{2F_{D}}{\rho U_{\perp}^{2}TL_{z}}\quad\text{and}\quad C_{L\perp}=\frac{2F_{L}}{\rho U_{\perp}^{2}CL_{z}}.$ (5)
As the whisker geometry lacks camber and is symmetric about the $xz$ plane, mean lift is zero and thus the magnitude of oscillations is captured by the root mean square (RMS) of the coefficient over $n$ samples in time with a sampling rate of 50 samples per convective time unit over the 400 CTU statistically developed period,
$C_{L,RMS}=\sqrt{\frac{1}{n}\sum_{i=1}C_{L_{i}}^{2}}.$ (6)
To provide a baseline comparison, simulations of flow over a circular cylinder and elliptical cylinder of equivalent mean thickness are performed. The elliptical cylinder maintains the same mean aspect ratio as the whisker model removing effects of the streamlined cross-section and serves as a useful intermediate comparison between a circular cylinder and the undulated cylinder.
Shedding frequency is computed from a fast Fourier transform (FFT) of the spanwise-averaged lift coefficient to quantify the dominant unsteady behavior across sweep angles. A reduced frequency $f^{*}$ is computed as
$f^{*}=\frac{fT}{U_{\infty}}.$ (7)
To aid in flow visualization $Q$-criterion is calculated to highlight vortices in the wake by relating the magnitude of local strain rate $\textbf{{S}}=\frac{1}{2}(\nabla\textbf{{u}}+\nabla\textbf{{u}}^{T})$ and rotation rate $\bm{\varOmega}=\frac{1}{2}(\nabla\textbf{{u}}-\nabla\textbf{{u}}^{T})$ tensors,
@@
Lastly, the distribution of energy in the wake is compared via contours of turbulent kinetic energy (TKE), defined by the half-sum of variances $\overline{(\textbf{{u}}^{\prime})^{2}}$ in fluctuating velocity $\textbf{{u}}^{\prime}$ obtained in the body-fixed coordinate system by subtracting time-averaged velocity $\overline{\textbf{{u}}}$ from instantaneous velocity $\textbf{{u}}(t)$,
$k=\frac{1}{2}\left[\overline{(u^{\prime})^{2}}+\overline{(v^{\prime})^{2}}+\overline{(w^{\prime})^{2}}\right].$ (9)